The equation of hyperbola with center at origin and co ordinate axes as its axes, distance between the directrices being 4/√3 and passing through the point (2,1) is?
pls explain in detail
thanks in advance

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
2a/e = 4/sqrt(3)..........(i)
Now
 
Hyperbola is of form
 
x²/a² - y²/b² = 1
 
It passes through (2,1)
 
4/a² - 1/b² = 1
 
4b² -a² = a² b²........(ii)
 
Now solve (i) and (ii)
 
You will get values of a² and b²
 
Hence you can get equation of hyperbola